# The Holomorphic Extension Property for $k$-du Bois Singularities

### v2

@@ 2.1. Notation
Throughout, we let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$. We will use $\pi:\tilde{X}\to X$ to denote a strong and projective resolution of singularities.
@@ 2.2. The du Bois Complex
Let $X$ be a complex algebraic variety. The du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$ is an object in the derived category of filtered complexes generalizing the holomorphic de Rham complex for algebraic varieties over $\mathbb{C}$. We denote its graded pieces by $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$. Studied by du Bois [DB81] from Deligne’s construction of the mixed Hodge structure [Del74], the du Bois complex is constructed by simplicial or cubical hyperresolutions of $X$. We will not need this construction in this paper but will only use its formal consequences. The interested reader can consult [PS08] for a good treatment of this construction.
@@
    [Kov11, Proposition 2.6], [SVV23, Lemma 3.2] If $H\subset X$ is a general member of a basepoint free linear system, then $\underline{\Omega}_{H}^{\bullet}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{\bullet}\otimes\mathscr{O}_{H}$.
@@
    where $\pi:\tilde{X}\to X$ is a morphism of complex schemes of finite type with singular locus $\Sigma\subset X$ and $E=\pi^{-1}(\Sigma)$. In particular, $\underline{\Omega}_{X}^{n}\cong_{\mathrm{qis}}\mathbf{R}\pi_{*}\omega_{\tilde{X}}\cong_{\mathrm{qis}}\pi_{*}\omega_{\tilde{X}}$ if $\pi:\tilde{X}\to X$ is a resolution of singularities.
@@
2.3. du Bois, $k$-du Bois, and $k$-rational Singularities
@@
    We say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
@@
    We say that $X$ has $k$-rational singularities if the natural map $\Omega_{X}^{p}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ is a quasi-isomorphism for each $0\leq p\leq k$.
@@
    We say that $X$ has weakly $k$-du Bois singularities, or pre $k$-du Bois singularities, if the natural map $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
@@
    We say that $X$ has weakly $k$-rational singularities, or pre $k$-rational singularities, if the natural map $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ for each $0\leq p\leq k$.
Here are some general properties of (weakly) $k$-du Bois (resp. (weakly) $k$-rational singularities).
  • By Theorem 2.1(vi), if $H$ is a general member of a basepoint linear system of a variety $X$ with (weakly) $k$-du Bois singularities, then $H$ also has (weakly) $k$-du Bois singularities (resp. (weakly) $k$-rational singularities).
@@
  • If $X$ is a normal variety and $f:Y\to X$ is a finite dominant map from a variety $Y$ with rational and weakly $k$-du Bois singularities, then $X$ also has rational [Kov00, Theorem 1] and weakly $k$-du Bois (resp. weakly $k$-rational) singularities [SVV23, Proposition 4.2].
  • If $X$ has simple normal crossing singularities, then $X$ is weakly $k$-du Bois since
@@
  • If $X$ has rational singularities, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{[p]}$ for each $p$, which follows from [KS21] and [HJ14]. In particular, a complex algebraic variety $X$ with rational singularities is weakly $k$-du Bois (resp. $k$-rational) if and only if $\Omega_{X}^{[p]}\to\underline{\Omega}_{X}^{p}$ (resp. $\Omega_{X}^{[p]}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$) is a quasi-isomorphism for each $0\leq p\leq k$.
There is a nice characterization due to Schwede which wish to mention. For $X$ a reduced and separated complex scheme of finite type, there exists a (local) embedding $\iota:X\to Y$ of $X$ into a smooth scheme $Y$. Let $\pi:\widetilde{Y}\to Y$ be an embedded resolution of $X$ which is an isomorphism outside of $X$, and let $\overline{X}=\pi^{-1}(X)_{\mathrm{red}}$ be the reduced preimage.
@@ 3.1.1. Pure and Mixed Hodge Modules
Let $Y$ be a smooth complex manifold of dimension $d$ (for example, let $Y$ be a local embedding of a complex algebraic variety $X$ of codimension $c=d-n$). A pure Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},\mathrm{rat}~{}M)$ consisting of:
@@
    and $\mathrm{gr}_{F}^{\bullet}$ is coherent $\mathrm{gr}_{F}^{\bullet}\mathscr{D}_{Y}$.
@@
A mixed Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)$ consisting of a $\mathscr{D}_{Y}$-module structure $\mathcal{M}$, a perverse structure $\mathrm{rat}~{}M$, a decreasing good filtration $F_{\bullet}$ on $\mathcal{M}$, and an increasing weight filtration $W_{\bullet}$ on these structures such that the graded pieces
$\mathrm{gr}_{k}^{W}M=(\mathrm{gr}_{k}^{W}\mathcal{M},\mathrm{rat}~{}\mathrm{gr}_{k}^{W}M,F_{\bullet})$
are pure Hodge modules. We call $M$ a graded-polarizable mixed Hodge module if further $\mathrm{gr}_{k}^{W}M$ is polarizable for each $k$. We refer to the category of graded-polarizable mixed Hodge modules as $\mathrm{MHM}(Y)$.
@@ 3.1.2. The Dual Mixed Hodge Module
Let $Y$ be a complex manifold of dimension $d$, and fix a mixed Hodge module $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
We define the dual Hodge module $\mathbb{D}_{Y}(M)$ to be the mixed Hodge module satisfying
(3.2) $\mathbb{D}_{Y}(\mathrm{gr}_{k}^{W}M)=\mathrm{gr}_{-k}^{W}\mathbb{D}_{Y}(M)$
for every $k$. The underlying perverse sheaf $\mathrm{rat}~{}\mathbb{D}(M)$ is the Verdier dual of $\mathrm{rat}~{}M$, and the underlying $\mathscr{D}_{Y}$-module is the holonomic dual
$\mathbb{D}_{Y}(\mathcal{M})=R^{d}\mathrm{Hom}_{\mathscr{D}_{Y}}(\omega_{Y}\otimes_{\mathscr{O}_{Y}}\mathcal{M},\mathscr{D}_{Y}),$
which is compatible with the filtration $F_{\bullet}$ of item (ii) of §3.1.1, by [Sai88, Lemma 5.1.13]. In particular, if $M$ is a pure Hodge module of weight $w$, then $\mathbb{D}_{Y}(M)$ is a pure Hodge module of weight $-w$. In this case, $\mathbb{D}_{Y}(M)\cong M(w)$, where $M(w)$ is the Tate-twist of $M$ in degree $w$, obtained by twisting the perverse structure and Hodge filtration in the usual way.
The compatibility of the Hodge filtration with $\mathbb{D}_{Y}$ is due to Saito [Sai88, 2.4.3]. Specifically, there is an isomorphism
(3.3) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{M})),$
where again $\omega_{Y}^{\bullet}=\omega_{Y}[d]$. If $M\in\mathrm{HM}(Y,w)$, the isomorphism $\mathbb{D}_{Y}(M)\cong M(w)$ reduces (3.3) to
@@ 3.1.3. The Restricted Hodge Module
Now let $Y=\mathbb{C}^{d}$ and $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$. For a generic hyperplane section $H$ of $Y$, $H$ intersects any Whitney strata adapted to the perverse sheaf $\mathrm{DR}(\mathcal{M})$ transversely. Therefore $H$ defines a non-characteristic hypersurface with respect to the left $\mathscr{D}_{Y}$-module $\mathcal{M}$, see [KS21, Definition 4.15] and [Sch16, §9]. Given such a hyperplane $H\subset Y$, we can construct a mixed Hodge module $M_{H}\in\mathrm{MHM}(H)$ by [Sai90, Lemma 2.25]. If $\iota_{H}:H\hookrightarrow Y$ is the inclusion, the underlying $\mathscr{D}_{H}$-module is
@@
with filtration $F_{\bullet}\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}F_{\bullet}\mathcal{M}$, see [Sch16, Lemma 9.5]. The de Rham complex of $\mathcal{M}_{H}$ is
@@ 3.2. The Intersection Hodge Module
We remark that the underlying perverse sheaf $IC_{X}$ of $\mathbb{Q}$-constructible sheaves computes the intersection cohomology groups $IH^{\bullet}(X,\mathbb{Q}):=\mathbb{H}^{\bullet-n}(X,IC_{X})$ of $X$. With this in mind, we get a Hodge-theoretic interpretation of the decomposition theorem.
Let $\pi:\widetilde{X}\to X$ be a resolution of singularities, and consider the induced morphism $f=\iota\circ\pi:\widetilde{X}\to Y$. By Saito’s direct image theorem [Sai88, §5.3], there are pure Hodge modules $M_{l}=(\mathcal{M}_{l},F_{\bullet},\mathrm{rat}~{}M_{l})$ supported on the singularities of $X$ and a decomposition
@@ 3.3. The Logarithmic Hodge Module
Let $X$ be a reduced and irreducible complex analytic variety and $\iota:X\hookrightarrow Y$ a (local) closed embedding into a complex ball $Y=\mathbb{C}^{n+c}$ of codimension $c$. If $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with log-exceptional divisor $E$, recall from §2.1 that we assume an isomorphism $U:=X_{\mathrm{reg}}\cong\widetilde{X}\setminus E$. The sheaf $\mathbb{Q}_{\widetilde{X}\setminus E}$ defines a variation of pure Hodge structures; as a perverse sheaf, we can consider the pushforward $\mathbf{R}j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}$, where $j:\tilde{X}\setminus E\hookrightarrow\widetilde{X}$ is the inclusion. By [Sai90, Thm. 3.27], this extends to a graded-polarizble mixed Hodge module on the smooth manifold $\widetilde{X}$. Specifically, the underlying $\mathscr{D}_{\widetilde{X}}$-module is the sheaf $\mathscr{O}_{\widetilde{X}}(*E)$ of meromorphic functions which are regular outside $E$. The filtered pieces $F_{p}\mathscr{D}_{\widetilde{X}}$ act naturally on $\mathscr{O}_{\widetilde{X}}(*E)$, inducing a good filtration on $\mathscr{O}_{\widetilde{X}}(*E)$. The de Rham complex is simply
@@
By [KS21, Proposition 9.5], we get the following important relationship between logarithmic forms and the logarithmic Hodge module $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$:

### v3

@@ 2.1. Notation
Throughout, we let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$. We will use $\pi:\tilde{X}\to X$ to denote a projective resolution of singularities.
@@ 2.2. The Du Bois Complex
Let $X$ be a complex algebraic variety. The Du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$ is an object in the derived category of filtered complexes of constructible sheaves $\mathscr{D}^{b}_{\mathrm{filt}}(X)$, generalizing the holomorphic de Rham complex for algebraic varieties over $\mathbb{C}$. We denote its graded pieces by $\underline{\Omega}_{X}^{p}:=\mathrm{gr}^{F}_{p}\underline{\Omega}_{X}^{\bullet}[p]$, which are defined in the bounded derived category of coherent sheaves. Studied by Du Bois [DB81] from Deligne’s construction of the mixed Hodge structure [Del74], the Du Bois complex is constructed by simplicial or cubical hyperresolutions of $X$. We will not need this construction in this paper but will only use its formal consequences. The interested reader can consult [PS08] for a good treatment of this construction.
@@
    [Kov11, Proposition 2.6], [SVV23, Lemma 3.2] If $H\subset X$ is a general member of a basepoint free linear system, then there is an exact triangle
    $\underline{\Omega}_{H}^{p-1}\otimes_{\mathscr{O}_{H}}^{\mathbb{L}}\mathscr{O}_{H}(-H)\to\underline{\Omega}_{X}^{p}\otimes_{\mathscr{O}_{X}}^{\mathbb{L}}\mathscr{O}_{H}\to\underline{\Omega}_{H}^{p}\xrightarrow{+1}.$
    In particular, $\underline{\Omega}_{X}^{0}\otimes\mathscr{O}_{H}\cong_{\mathrm{qis}}\underline{\Omega}_{H}^{0}$.
@@
    where $\pi:\tilde{X}\to X$ is a resolution of singularities, $\Sigma\subset X$ is the singular locus, and $E=\pi^{-1}(\Sigma)$. In particular, $\underline{\Omega}_{X}^{n}\cong_{\mathrm{qis}}\mathbf{R}\pi_{*}\omega_{\tilde{X}}\cong_{\mathrm{qis}}\pi_{*}\omega_{\tilde{X}}$.
@@
2.3. Du Bois, $m$-Du Bois, and $m$-rational Singularities
@@
    We say that $X$ has $m$-Du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq m$.
@@
    We say that $X$ has $m$-rational singularities if the natural map $\Omega_{X}^{p}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ is a quasi-isomorphism for each $0\leq p\leq m$.
@@
    We say that $X$ has weakly $m$-Du Bois singularities if the natural map $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq m$.
@@
    We say that $X$ has weakly $m$-rational singularities if the natural map $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ for each $0\leq p\leq m$.
Here are some general properties of (weakly) $m$-Du Bois (resp. (weakly) $m$-rational singularities).
  • If $H$ is a general member of a basepoint linear system of a variety $X$ with (weakly) $m$-Du Bois singularities, then $H$ also has (weakly) $m$-Du Bois singularities (resp. (weakly) $m$-rational singularities) [SVV23, Theorem A, Corollary 3.3].
@@
  • If $X$ is a normal variety and $f:Y\to X$ is a finite dominant map from a variety $Y$ with rational and weakly $m$-Du Bois singularities, then $X$ also has rational [Kov00, Theorem 1] and weakly $m$-Du Bois (resp. weakly $m$-rational) singularities [SVV23, Proposition 4.2].
  • If $X$ has simple normal crossing singularities, then $X$ is weakly $m$-Du Bois since
@@
  • If $X$ has rational singularities, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{[p]}$ for each $p$, which follows from [KS21] and [HJ14]. In particular, a complex algebraic variety $X$ with rational singularities is weakly $m$-Du Bois (resp. $m$-rational) if and only if $\Omega_{X}^{[p]}\to\underline{\Omega}_{X}^{p}$ (resp. $\Omega_{X}^{[p]}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$) is a quasi-isomorphism for each $0\leq p\leq k$.
There is a nice characterization due to Schwede. For $X$ a reduced and separated complex scheme of finite type, there exists a (local) embedding $\iota:X\to Y$ of $X$ into a smooth scheme $Y$. Let $\pi:\widetilde{Y}\to Y$ be an embedded resolution of $X$ which is an isomorphism outside of $X$, and let $\overline{X}=\pi^{-1}(X)_{\mathrm{red}}$ be the reduced preimage.
@@ 3.1.1. Pure and Mixed Hodge Modules
Let $Y$ be a smooth complex manifold of dimension $d$ (for example, let $Y$ be a local embedding of a complex algebraic variety $X$ of codimension $c=d-n$). A pure Hodge module on $Y$ is an object $M=(\mathcal{M},F^{\bullet},\mathrm{rat}~{}M)$ consisting of:
@@
    and $\mathrm{gr}^{F}_{\bullet}\mathcal{M}$ is coherent over $\mathrm{gr}^{F}_{\bullet}\mathscr{D}_{Y}$.
@@
A mixed Hodge module on $Y$ is an object $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)$ consisting of a $\mathscr{D}_{Y}$-module structure $\mathcal{M}$, a perverse structure $\mathrm{rat}~{}M$, a decreasing good filtration $F^{\bullet}$ on $\mathcal{M}$, and an increasing weight filtration $W_{\bullet}$ on these structures such that the graded pieces
$\mathrm{gr}_{W}^{k}M=(\mathrm{gr}_{W}^{k}\mathcal{M},\mathrm{rat}~{}\mathrm{gr}_{W}^{k}M,F^{\bullet})$
are pure Hodge modules. We call $M$ a graded-polarizable mixed Hodge module if further $\mathrm{gr}_{W}^{k}M$ is polarizable for each $k$. We refer to the category of graded-polarizable mixed Hodge modules as $\mathrm{MHM}(Y)$.
@@ 3.1.2. The Dual Mixed Hodge Module
Let $Y$ be a complex manifold of dimension $d$, and fix a mixed Hodge module $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$. There is a mixed Hodge module $\mathbf{D}_{Y}(M)$, known as the dual Hodge module, satisfying
(3.2) $\mathbf{D}_{Y}(\mathrm{gr}_{W}^{k}M)=\mathrm{gr}_{-k}^{W}\mathbf{D}_{Y}(M)$
for every $k$. The underlying perverse sheaf $\mathrm{rat}~{}\mathbf{D}_{Y}(M)$ is the Verdier dual of $\mathrm{rat}~{}M$, and the underlying $\mathscr{D}_{Y}$-module is the holonomic dual
$\mathbf{D}_{Y}(\mathcal{M})=R^{d}\mathrm{Hom}_{\mathscr{D}_{Y}}(\omega_{Y}\otimes_{\mathscr{O}_{Y}}\mathcal{M},\mathscr{D}_{Y}),$
which is compatible with the filtration $F^{\bullet}$ of item (ii) of §3.1.1, by [Sai88, Lemma 5.1.13]. In particular, if $M$ is a pure Hodge module of weight $w$, then $\mathbf{D}_{Y}(M)$ is a pure Hodge module of weight $-w$. In this case, $\mathbf{D}_{Y}(M)\cong M(w)$, where $M(w)$ is the Tate-twist of $M$ in degree $w$, obtained by twisting the perverse structure and Hodge filtration in the usual way.
The compatibility of the Hodge filtration with $\mathbf{D}_{Y}$ is due to Saito [Sai88, 2.4.3]. Specifically, there is an isomorphism
(3.3) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{M})),$
where again $\omega_{Y}^{\bullet}=\omega_{Y}[d]$. If $M\in\mathrm{HM}(Y,w)$, the isomorphism $\mathbf{D}_{Y}(M)\cong M(w)$ reduces (3.3) to
@@ 3.1.3. The Restricted Hodge Module
Now let $Y=\mathbb{C}^{d}$ and $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$. For a generic hyperplane section $H$ of $Y$, $H$ intersects any Whitney strata adapted to the perverse sheaf $\mathrm{DR}(\mathcal{M})$ transversely. Therefore $H$ defines a non-characteristic hypersurface with respect to the left $\mathscr{D}_{Y}$-module $\mathcal{M}$, see [KS21, Definition 4.15] and [Sch16, §9]. Given such a hyperplane $H\subset Y$, we can construct a mixed Hodge module $M_{H}\in\mathrm{MHM}(H)$ by [Sai90, Lemma 2.25]. If $\iota_{H}:H\hookrightarrow Y$ is the inclusion, the underlying $\mathscr{D}_{H}$-module is
@@
with filtration $F^{\bullet}\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}F^{\bullet}\mathcal{M}$, see [Sch16, Lemma 9.5]. The de Rham complex of $\mathcal{M}_{H}$ is
@@ 3.2. The Intersection Hodge Module
We remark that the underlying perverse sheaf $IC_{X}$ computes the intersection cohomology $IH^{\bullet}(X,\mathbb{Q}):=\mathbb{H}^{\bullet-n}(X,IC_{X})$ of $X$. With this in mind, we get a Hodge-theoretic interpretation of the decomposition theorem.
Let $\pi:\widetilde{X}\to X$ be a resolution of singularities, and consider the induced morphism $f=\iota\circ\pi:\widetilde{X}\to Y$. By Saito’s direct image theorem [Sai88, §5.3], there are pure Hodge modules $M_{l}=(\mathcal{M}_{l},F^{\bullet},\mathrm{rat}~{}M_{l})$ supported on the singularities of $X$ and a decomposition
@@ 3.3. The Logarithmic Hodge Module
Let $X$ be a reduced and irreducible complex analytic variety and $\iota:X\hookrightarrow Y$ a (local) closed embedding into a complex ball $Y=\mathbb{C}^{n+c}$ of codimension $c$. If $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with log-exceptional divisor $E$, recall from §2.1 that we assume an isomorphism $U:=X_{\mathrm{reg}}\cong\widetilde{X}\setminus E$. The sheaf $\mathbb{Q}_{\widetilde{X}\setminus E}[n]$ defines a variation of pure Hodge structures; as a perverse sheaf, we can consider the pushforward $\mathbf{R}j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}$, where $j:\tilde{X}\setminus E\hookrightarrow\widetilde{X}$ is the inclusion. By [Sai90, Thm. 3.27], this extends to a graded-polarizble mixed Hodge module on the smooth manifold $\widetilde{X}$. Specifically, the underlying $\mathscr{D}_{\widetilde{X}}$-module is the sheaf $\mathscr{O}_{\widetilde{X}}(*E)$ of meromorphic functions which are regular outside $E$. The filtered pieces $F_{p}\mathscr{D}_{\widetilde{X}}$ act naturally on $\mathscr{O}_{\widetilde{X}}(*E)$, inducing a good filtration on $\mathscr{O}_{\widetilde{X}}(*E)$. The de Rham complex is simply
@@
By [KS21, Proposition 9.5], we get the following important relationship between logarithmic forms and the logarithmic Hodge module $N_{0}=(\mathcal{N}_{0},F^{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$: